OpenStudy (anonymous):

Solve the following question: (-sin^2)x=2cosx-2

OpenStudy (luigi0210):

Start by moving everything to one side

OpenStudy (luigi0210):

\[-\sin^2x-2cosx+2=0\] \[-(1-\cos^2x)-2cosx+2=0\] \[\cos^2x-1-2cosx+2=0\] \[\cos^2x-2cosx+1=0\] \[(cosx-1)(cosx-1)=0\]

OpenStudy (anonymous):

okay so would that be the final answer? im kinda confused I dont know if should be proofing it or solving for x

OpenStudy (luigi0210):

no, the final answer is: \[cosx=\pm1\] where cos equals +/-1

OpenStudy (luigi0210):

so what trig values equal +/-1?

OpenStudy (anonymous):

would you do x=(cos^-1)(1)

OpenStudy (anonymous):

you would find the inverse

OpenStudy (luigi0210):

You could.. you could also use the unit circle and see where cos is 1/-1